1. The switch S is closed at t = 0 (assume that the battery voltage remains constant at 10V and the resistance of the inductor is negligible). Calculate the voltage across each resistor a very long time after the switch has been closed and all currents and voltages reached steady values. (5 points)1. The switch S is closed at t = 0 (assume that the battery voltage remains constant at 10V and the resistance of the inductor is negligible). Calculate the voltage across each resistor a very long time after the switch has been closed and all currents and voltages reached steady values. (5 points)10 22 1. The switch S is closed at t = 0 (assume that the battery voltage remains constant at 10V and the resistance of the i
Answer #1

1)

A long time after the switch is closed, the current through the inductance does not change . Voltage across the inductor is V_{2H}=L\frac{dI}{dt}=0 . It can be considered as zero resistance.It is in parallel with 100\,\Omega resistor, so resistance of the combination will be zero. Now the circuit consists of a battery of emf OV and a resistance 10\,\Omega .

Current through the circuit is I=\frac{10}{10}=1\,A

Voltage across the 10\,\Omega resistor is V_{10\,\Omega}=10\,V

Voltage across the inductor is V_{2H}=0\,V

Voltage across the 100\,\Omega resistor is V_{100\,\Omega}=0\,V

Current through 10\,\Omega resistor is I_{10\,\Omega}=1\,A

Current through inductor is I_{2H}=1\,A

Current through 100\,\Omega resistor is I_{100\,\Omega}=0\,A

2)

After currents and voltages reached steady state again, the switch is opened again.

Now the circuit consists of 100\,\Omega resistor and 2H inductor in series. At time t=0 , the current through the circuit is I_0=1\,A ,

At time t=0 , the current through the  2H inductor is I_{2H}=1\,A and is flowing down the inductor.

Current through the 100\,\Omega resistor after time t is I_{100\,\Omega}=I_0(1-e^{-Rt/L})

I_{100\,\Omega}=1(1-e^{-100*0.02/2})=0.632\,A

Direction of current in 100\,\Omega resistor is UP.